Assume that a five percent (5%) increase in the price of a product leads to a ten percent (10%) drop in the quantity demanded for the product. Describe in full what type of good this will be defined as, according to the theory of price elasticity of demand. Motivate your answer.
1
Expert's answer
2019-09-10T13:16:49-0400

Е = ∆Q / ∆P (%)

E - elasticity

∆Q = (q2-q1)/q1

∆P = (p2-p1)/p1

q - quantity

p - price

∆Q = (0.9q1-q1)/q1=-0.1q1/q1 = -0.1

∆P = (1.05p1-p1)/p1 = 0.05p1/p1 = 0.05

E = -0.1/0.05 = -2

We see that the value of elasticity is greater than 1. In this case, the elasticity value is determined modulo.

Products with elastic demand at a price:

 * Luxury goods (jewelry, delicacies)

 * Goods whose value is tangible for the family budget (furniture, household appliances)

 * Easily replaceable goods (tea, coffe, meat, fish, fruits, vegetables)